Find the p-value for the hypothesis test. A random sample of size 51 is taken. The sample has a mean of 392 and a standard deviation of 83.

H0: µ = 400

Ha: µ< 400

The p-value for the hypothesis test is .

Your answer should be rounded to 4 decimal places.

Question 16
Answer #1

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 400

Ha : \mu < 400

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (392-400) / 83 / \sqrt 51

= -0.69

P(z <-0.69 ) = 0.2451

P-value = 0.2451